# On Nontrivial Winning and Losing Parameters of Schmidt Games

### v1

@@ Abstract.
In this paper we completely describe the winning and losing conditions different from the only “trivial” conditions known before. In other words, we solve the open question of finding a complete nontrivial Schmidt diagram. In addition, we give the new bounds for two family of sets: one related to frequencies of digits in base-$2$ expansions, and one connected to the set of the badly approximable numbers.
@@ 1. Set-up and Introduction: Schmidt Games
Wolfgang Schmidt first introduced Schmidt games in [S66] in order to study the set of badly approximable numbers. Winning sets in these games have special properties, such as a full Hausdorff dimension; as such, Schmidt games have been used to prove results in number theory and dynamical systems. Here, we briefly describe the setup of a Schmidt $(\alpha,\beta)$-game. Throughout this paper we work with the Schmidt game on $\mathbb{R}$, so for simplicity we give the definitions only in this special case.
@@
Then, fix a set $S\subset{\mathbb{R}}$ and $(\alpha,\beta)\in I$. Two players, Alice and Bob, can play a Schmidt game on $S$, taking alternating turns, as follows: Bob begins by choosing an interval $B_{0}$. Given an interval $B_{i}$, Alice chooses an interval $A_{i}\subset B_{i}$ such that $|A_{i}|=\alpha|B_{i}|$. Similarly, given an interval $A_{i}$, Bob chooses an interval $B_{i+1}\subset A_{i}$ such that $|B_{i+1}|=\beta|A_{i}|$. Let
@@
We call $S$ $(\alpha,\beta)$-winning if, regardless of how Bob chooses his intervals, Alice can play in such a way that $x_{\infty}\in S$. For every $S$, define the Schmidt diagram $D(S)$ of $S$ as the set of all pairs $(\alpha,\beta)\in I$ such that $S$ is $(\alpha,\beta)$-winning. In his original paper, Schmidt proved that there are two “trivial zones” of Schmidt diagrams:
@@
There are only four sets which were known to coincide with $D(S)$ for some $S$: $\emptyset$, $\check{D}$, $\hat{D}$, and $I$. Our goal with this project was to find some set $S$ such that $\check{D}\subsetneq D(S)\subsetneq\hat{D}$ and to completely describe $D(S)$. It is worth noting that this problem was investigated in an earlier PRIMES paper, which investigated sets relating to particular decimal expansions [KZT12]. In the aforementioned paper, some nontrivial statements about Schmidt diagrams were obtained, but no $D(S)$ different from the four diagrams discussed above was completely described.
@@ 2. Main Results
One of the main results of this paper is a complete description of a Schmidt Diagram different from $\emptyset$, $\check{D}$, $\hat{D}$ and $I$. More specifically, we prove the following theorem.
@@
There exists a set $\mathscr{S}$ such that $D(\mathscr{S})=\{(\alpha,\beta):0<\alpha<\beta<1\}$ (see Figure 3).
Figure 3. $D(\mathscr{S})$ (blue)
We develop the techniques for this in Section 3, and the more general statements about sets with such a diagram are given in Theorem 3.10 and Corollary 3.11. We construct a special family of examples in Section 4.
In Section 5, we describe some winning and losing conditions for families of sets connected with digit frequencies in the dyadic expansion of real numbers. We prove a general case in Theorem 5.3 and a special symmetric case in Corollary 5.6. In Section 6 we give some new results about winning and losing conditions for families of sets connected with badly approximable numbers in base $B$ expansion: specifically in Theorems 6.4, 6.5, and 6.6.
The bounds for the winning and losing zones we describe in Sections 5 and 6 are not optimal; however, they give some estimates and shed some light on how their respective Schmidt diagrams can look like. Figures 5 and 5 show two examples of the possible bounds (where the blue zones are subsets of the corresponding Schmidt diagram and the red zones are subsets of its complement). The sets $D_{c}^{-}$ used in the first example will be defined later in Section 5 and the sets $\operatorname{\widetilde{2-BA}}(c)$ will be defined in Section 6.
Figure 4. Bounds for $D(D_{c}^{-})$ with $c=0.7$
Figure 4. Bounds for $D(D_{c}^{-})$ with $c=0.7$
@@
3. General strategies in the Schmidt’s game
This section gives some general statements concerning strategies in Schmidt’s game.
@@ 3. General strategies in the Schmidt’s game
Fix $m\in\mathbb{N}$. Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$. Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $D$ that allows him to win, and the first ball in this strategy is $B_{0}$. Then:
@@
    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $D$ starting with the ball $m^{k}B_{0}$;
@@
    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $D$ starting with the ball $m^{k}+B_{0}$.
@@
    It is easy to see that $\bigcap\limits_{i=0}^{\infty}B_{i}^{\prime}=m^{k}\bigcap\limits_{i=0}^{\infty}B_{i}=m^{k}x$, where $x$ is the outcome of some game with the starting ball $B_{0}$, i.e. $x\notin D$, therefore $m^{k}x\notin D$ and Bob wins.
@@
    The strategy and the argument is analogous; we just take $A_{i}=A_{i}^{\prime}-m^{k}$ and $B_{i}^{\prime}=B_{i}+m^{k}$ and use the fact that $x\in D$ if and only if $m^{k}+x\in D$.
@@
Fix $m\in\mathbb{N}$. Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$. Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $D$. Then for any $L>0$ and any $\varepsilon>0$ there exists a winning strategy in this game for Bob starting with $B_{0}$ with
@@
Fix $m\in\mathbb{N}$. Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$. Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $D$. Then for any $\beta^{\prime}<\beta,\,\alpha^{\prime}=\frac{\alpha\beta}{\beta^{\prime}}$, and any closed ball $T$ Bob can win $(\alpha^{\prime},\beta^{\prime})$-game starting with any ball $T$ as his initial ball.
@@
The discussion above implies the following statement:
@@
Let $m\in\mathbb{N}$ and $D,Q\subseteq\mathbb{R}$ be some sets with the following properties:
  • $m^{k}D=m^{k}+D=D$ for any integer $k$.
  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$.
@@
    If $\alpha\geq\beta$, then the set $D$ is $(\alpha,\beta)$-losing.
@@
    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $D\sqcup Q$ is $(\alpha,\beta)$-winning.
@@
    Suppose $\alpha\geq\beta$ and $D$ is $(\alpha,\beta)$-winning. In this case Alice has a positional winning strategy described by the function $A(I)$ (where $I\subset\mathbb{R}$ is a closed interval) for this game. Consider $(\beta,\alpha)$-game on the set $D$ and let Bob play the following strategy:
@@
    Suppose $x=\bigcap\limits_{i=0}^{\infty}B_{i}$ is the outcome of the game. Then $-x=\bigcap\limits_{i=0}^{\infty}A(-A_{n})$ is the outcome of the $(\alpha,\beta)$-game where Bob plays the balls $[-1;0];-A_{0},-A_{1},\ldots$ and Alice plays using her winning strategy $A(I)$. Therefore, $-x\in D$, which means $-x\notin D$ and Bob wins.
    It means that the point $(\alpha,\beta)$ is winning and the point $(\beta,\alpha)$ is losing for the set $D$, which contradicts Lemma 3.1. Thus, the point $(\alpha,\beta)$ is losing.
@@
    The idea of the proof is the same as in (a). Suppose $D\sqcup Q$ is $(\alpha,\beta)$-losing; Take $\varepsilon>0$ under the condition $(1+\varepsilon)^{2}<\frac{\beta}{\alpha}$. Then by Lemma 3.7 for any closed ball $T$ Bob has a winning strategy for $\left(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)}\right)$-game starting from $T$. We use this statement to construct a $\left(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon)\right)$-winning strategy for Alice.
    Suppose Bob starts some $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$-game with the ball $B_{0}$. Then Alice picks some ball $A_{0}\subseteq B_{0}$ without any restrictions. If Bob plays the ball $B_{n}$ on $n$-th step, Alice responds by playing $A_{n}$ where $-A_{n}$ is the ball Bob would play in $(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)})$-game which started from $-A_{0}$ after the balls $-A_{0},-B_{1},\ldots,-B_{n}$ (in this game $-A_{i}$ are Bob’s balls and $-B_{i}$ are Alice’s). Let $x$ be the outcome of this $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$-game; then $-x$ is the outcome of $(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)})$-game Bob played with his winning strategy, i.e. $-x\in(-D)$ and $x\in D\subseteq D\sqcup Q$. It shows that $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$ is a winning point for $D\sqcup Q$.
    But one has $\alpha<\alpha(1+\varepsilon))<\frac{\beta}{(1+\varepsilon)}$, which by Lemma 3.1 contradicts our initial assumption. Therefore $D\sqcup Q$ is $(\alpha,\beta)$-winning.
Now, the goal of the next part is to find the condition under which we can remove dyadic rationals from the set $D$ so that it remains winning for the parameters above.
We start with the following almost obvious statement:
@@
Suppose $S$ is an $(\alpha,\beta)$-winning set and for some $\xi\in\mathbb{R}$ one has $S+\xi\subseteq D$. Then the set $D$ is $(\alpha,\beta)$-winning.

### v3

@@ Abstract.
In this paper we study the classical Schmidt game on two families of sets: one related to frequencies of digits in base-$2$ expansions, and one connected to the set of the badly approximable numbers. Namely, we describe some nontrivial winning and losing parameters $(\alpha,\beta)$ for these sets.
@@ 1. Set-up and Introduction: Schmidt Games
Wolfgang Schmidt first introduced Schmidt games in [S66] in order to study the set of badly approximable numbers. These games involve two players, Alice and Bob and two parameters $\alpha,\beta\in(0,1)$. A subset $S$ of a metric space is called $(\alpha,\beta)$-winning if Alice has a strategy guaranteeing that a certain point, constructed as a result of the infinite series of moves, lies in $S$. Of particular interest are sets that are $(\alpha,\beta)$-winning for all $\beta$ and a fixed $\alpha$; those are called $\alpha$-winning sets. Schmidt proved that $\alpha$-winning sets have full Hausdorff dimension, and a countable intersection of $\alpha$-winning sets is $\alpha$-winning. The technique of Schmidt games has been frequently used to prove that certain sets arising in number theory and dynamical systems have full Hausdorff dimension. Many of the applications of Schmidt games deal with various stronger modifications of Schmidt’s original construction. These modifications include the McMullen’s strong and absolute winning games [M10], with an additional property that the winning property is preserved by quasisymmetric homeomorphisms; hyperplane absolute winning game [BFKRW12], where the winning property holds under $C^{1}$ diffeomorphisms; potential game [FSU18], Cantor game [BH17], and Hausdorff/packing games [DFSU24] which give a tool to precisely calculate Hausdorff/packing dimensions. For a detailed survey describing the setups of some of the aforementioned games and connections between them, see [BHNS25]. However, in this paper, we restrict our attention to the $(\alpha,\beta)$-game introduced in [S66].
Here, we briefly describe the setup of the classical Schmidt $(\alpha,\beta)$-game. Throughout this paper we work with the Schmidt game on $\mathbb{R}$, so for simplicity we give the definitions only in this special case.
@@
Then, fix a set $S\subset{\mathbb{R}}$ and $(\alpha,\beta)\in I$. Two players, Alice and Bob, can play a Schmidt game on $S$, taking alternating turns, as follows: Bob begins by choosing a closed interval $B_{0}$. Given an interval $B_{i}$, Alice chooses a closed interval $A_{i}\subset B_{i}$ such that $|A_{i}|=\alpha|B_{i}|$. Similarly, given an interval $A_{i}$, Bob chooses an interval $B_{i+1}\subset A_{i}$ such that $|B_{i+1}|=\beta|A_{i}|$. Let
@@
We call $S$ $(\alpha,\beta)$-winning if, regardless of how Bob chooses his intervals, Alice can play in such a way that $x_{\infty}\in S$.
In this paper, we will often fix the set $S$ and explore the games for different pairs $(\alpha,\beta)$. To do this, we will utilize a little bit different language.
Let’s fix the set $S$. A point $(\alpha,\beta)\in I$ is called winning for $S$ if $S$ is $(\alpha,\beta)$-winning and non-winning for $S$ otherwise. A point $(\alpha,\beta)\in I$ is called losing for $S$ if, regardless of how Alice chooses her intervals, Bob can play in such a way that $x_{\infty}\notin S$, and non-losing otherwise.
It will be crucial for some of our proofs to know that for a given set $S$ and any parameters $(\alpha,\beta)$ one of the players always has a strategy to win; that is, we will require that all the non-losing parameters are winning, and all the non-winning parameters are losing. In this case, we will say that the Schmidt game is determined on the set $S$. In general, Schmidt game can be not determined: see [AFG23] for some constructions using the Axiom of choice. However, it is known ([CFJ23] ) that the Schmidt game in $\mathbb{R}$ is determined if we assume the Axiom of determinacy AD (which contradicts the Axiom of choice). More specifically, it is known that the Schmidt game is determined on Borel sets.
For every $S$, define the Schmidt diagram $D(S)$ of $S$ as the set of all winning $(\alpha,\beta)\in I$. In his original paper, Schmidt proved that there are two “trivial zones” of Schmidt diagrams:
@@
There are only four sets that are known to coincide with $D(S)$ for some $S$: $\emptyset$, $\check{D}$, $\hat{D}$, and $I$. It is clear that this is far not the complete list of the possible Schmidt diagrams, see [F82] for an example of the set which is proven to have a nontrivial diagram, however, no diagram $D(S)$ different from the four diagrams discussed above was completely described. It is worth noting that the problem of completely describing a nontrivial Schmidt diagram was investigated in an earlier PRIMES paper [KZT12], which investigated sets relating to particular decimal expansions.
In this paper, we partially describe the nontrivial Schmidt diagrams of two families of sets, and make some conjectures about their precise form.
The structure of the paper is as follows. In Section 2, we formulate our main results. We describe some winning and losing conditions for families of sets connected with digit frequencies in the dyadic expansion of real numbers in Subsection 2.1, and for families of sets connected with badly approximable numbers in base $B$ expansion in Subsection 2.2. In Section 3, we discuss some general properties of winning sets and general symmetric strategies on special families of sets, which we need to prove Theorem 2.2. We prove Theorem 2.2 in Section 4 and the theorems from Subsection 2.2 in Section 5.
@@ 2. Main results
2.1. Digit frequencies and related sets
We will investigate a construction regarding frequencies of digits in base-$2$ expansions of real numbers. These sets have been studied in the past (see [V59, FFNS21]).
First, some setup. Consider the base-$2$ expansions of the form $x=x_{0},x_{1}x_{2}\cdots$ where $x_{0}$ is an integer and $x_{i}\in\{0,1\}$ are the digits in the base-$2$ expansion of $x$. In this paper, if $x$ can be represented by two separate expansions, we take the lexicographically minimal expansion (for example, $0.100\ldots$ would instead be written as $0.011\ldots$). We define
and similarly,
Define the sets
$F^{+}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{+}(x,0)>c\},$
$F^{-}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{-}(x,0)>c\}.$
It is known ([Bo09]) that the set $\{x\in\mathbb{R}:\,\,d^{-}(x,j)=d^{+}(x,j)=\frac{1}{2}\}$ has full Lebesgue measure for any $j\in\{0,1\}$. In particular, the sets $F^{\pm}_{c}$ have full Lebesgue measure if $c<\frac{1}{2}$ and measure zero otherwise. Moreover, Besicovitch ([B35]; see also [O04] for generalizations) showed that the Hausdorff dimension of $F^{-}_{c}$ for $c>1$ is equal to $-c\log_{2}c-(1-c)\log_{2}(1-c)$.
Winning and losing parameters of these sets were studied before. The following fact is not stated, but follows from the proof of Theorem 2.15 in [FFNS21].
Proposition 2.1.
[FFNS21] Let $c\in(0,1)$. The set $F^{+}_{c}$ is winning for $\log(8\alpha)/\log(\alpha\beta)>c$ and losing for $\log(\alpha/8)/\log(\alpha\beta)\leq c$.
The results of our research expand on that bound:
@@ 2.1. Digit frequencies and related sets
Let $c\in(0,1)$, and suppose $\log_{2}{\alpha\beta}\notin\mathbb{Q}$. Then, the sets $F^{\pm}_{c}$ are $(\alpha,\beta)$-winning if $\log(4\alpha)/\log(\alpha\beta)>c$, and $(\alpha,\beta)$-losing if $\log(\alpha/4)/\log(\alpha\beta)\leq c$. In addition:
  1. (1)
    If $c<\frac{1}{2}$, then the set $F^{+}_{c}$ is $(\alpha,\beta)$-winning for any $(\alpha,\beta)$ such that $\beta>\alpha$ and $\log_{2}(\alpha\beta)\notin\mathbb{Q}$.
  2. (2)
    If $c\geq\frac{1}{2}$, then the set $F^{-}_{c}$ is $(\alpha,\beta)$-losing for any $(\alpha,\beta)$ such that $alpha\geq\beta$. Here we do not require the restriction $\log_{2}{\alpha\beta}\notin\mathbb{Q}$.
Let us also notice that Schmidt’s original approach (see Corollary 1 of Theorem 6 in [S66]) automatically gives us some losing zone bounds as a function of Hausdorff dimension. Namely, for any set $S\subseteq\mathbb{R}$ the following statement holds: if $\dim_{H}S=s$, and $\frac{\log\lfloor\beta^{-1}\rfloor}{-\log(\alpha\beta)}>s$, then $S$ is $(\alpha,\beta)$-non-winning. Using the fact that $\dim_{H}F_{c}^{-}=-c\log_{2}c-(1-c)\log_{2}(1-c)$ for $c>\frac{1}{2}$, we automatically get the following statement:
Proposition 2.3.
Let $c>\frac{1}{2}$, and suppose $\frac{\log\lfloor\beta^{-1}\rfloor}{-\log(\alpha\beta)}>-c\log_{2}c-(1-c)\log_{2}(1-c)$. Then, the set $F_{c}^{-}$ is $(\alpha,\beta)$-losing.
The winning and losing zones for selected values of $c$ are shown at figures  4,  4,  6 and  6. The winning zones from Theorem 2.2 are in blue, the losing zones from Theorem 2.2 are in red and the losing zones from Proposition 2.3 are in green. We note that the losing conditions in Theorem 2.2 are always stronger that the ones from Proposition 2.3 for small enough $\alpha$.
Figure 3. Bounds for $D(F^{+}_{c})$ with $c=0.1$
Figure 4. Bounds for $D(F^{+}_{c})$ with $c=0.55$
Figure 5. Bounds for $D(F^{-}_{c})$ with $c=0.9$
Figure 6. Bounds for $D(F^{-}_{c})$ with $c=0.45$
We will develop the necessary techniques and prove Theorem 2.2 in Section 3. Here, we would like to state a conjecture which our methods and results lead to.
Conjecture 2.4.
We believe that $D(F^{\pm}_{c})=\{(\alpha,\beta):\,\log(\alpha)/\log(\alpha\beta)>c\}$. In particular, we expect that $D(F^{\pm}_{\frac{1}{2}})=\{(\alpha,\beta):\,\alpha<\beta\}$.
2.2. d-Badly approximable numbers
It is known that the badly approximable numbers are dense and have Lebesgue measure zero. Schmidt showed in [S66] that they are winning for the largest possible bounds: namely, the set $\operatorname{BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$. That is, $D(\operatorname{BA})=\hat{D}$; in particular, the set $\operatorname{BA}$ is $\alpha$-winning for $\alpha<\frac{1}{2}$ and $\dim_{H}\operatorname{BA}=1$.
In a similar vein to the badly approximable numbers investigated by Schmidt, we define the set of d-Badly Approximable Numbers as
$\operatorname{d-BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{d^{k}}\right|>\frac{c}{d^{k}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}$
and a closely related set
$\operatorname{d-BA}(c):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{d^{k}}\right|>\frac{c}{d^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
These definitions imply $\operatorname{d-BA}=\bigcup_{c>0}\operatorname{d-BA}(c)$.
It also follows from the work of Schmidt ([S66]) that the set $\operatorname{2-BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{2-BA})=\hat{D}$.
While both $\operatorname{2-BA}$ and $\operatorname{2-BA}(c)$ did not produce any new Schmidt diagrams, there is a simple change we can make that will produce both nontrivial winning and nontrivial losing zones. For any integer $d\geq 2$, define
$\operatorname{\widetilde{d-BA}}(c,N):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{d^{k}}\right|>\frac{c}{d^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\ \text{s.t.}\ k>N\right\}$
and
$\operatorname{\widetilde{d-BA}}(c):=\bigcup_{N\in{\mathbb{N}}}\operatorname{\widetilde{d-BA}}(c,N).$
We note that these sets has been previously studied in [N10] and [PS08]. In particular, the Lebesgue measure of all the aforementioned sets is zero, and the precise Hausdorff dimension of $\operatorname{2-BA}$ for any parameter $c$ is calculated in [N10]. Moreover, there exists such a critical value $c_{0}$ that the set $\operatorname{2-BA}$ is countable if and only if $c>c_{0}$. This value is known to be related to Thue-Morse sequences; see Theorem 25 in [N10] for details.
Let us define $\gamma=1-2\alpha+\alpha\beta$.
The following theorems provide the winning and losing conditions for $\operatorname{\widetilde{d-BA}}$ sets.
Theorem 2.5.
Fix $0<c<\frac{1}{2}$. If
$\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma}>1,$
then for any $d$, the set $\operatorname{\widetilde{d-BA}}(c)$ is $(\alpha,\beta)$-losing.
Theorem 2.6.
Fix $0<c<\frac{1}{6}$. The set $\operatorname{\widetilde{2-BA}}(c)$ is $(\alpha,\beta)$-winning if $\alpha<\frac{1}{3}$ and $\beta>\frac{9c}{1-3\alpha}$.
Just as with the sets $F_{c}^{-}$, we note that some losing parameters for the sets $\operatorname{\widetilde{2-BA}}$ can be found using the Hausdorff dimension of $\operatorname{\widetilde{2-BA}}$. It is not hard to see that for small enough $c$ these losing zones we can obtain this way are contained in the zones given by Theorem 2.5.
A special case of winning and losing zones given by Theorem 2.5 and Theorem 2.6 can be seen at Figure 7.
@@
3. General strategies in the Schmidt game
This section gives some general statements concerning strategies in Schmidt’s game. These statements will be used to prove Theorem 2.2; however, we believe they may also prove useful for proving results concerning Schmodt games on a larger variety of sets.
@@ 3. General strategies in the Schmidt game
Fix $m\in\mathbb{N}$. Let $S\subseteq\mathbb{R}$ be some set with a property that $m^{k}=m^{k}+S=S$ for any integer $k$. Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $S$ that allows him to win, and the first ball in this strategy is $B_{0}$. Then:
@@
    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $S$ starting with the ball $m^{k}B_{0}$;
@@
    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $S$ starting with the ball $m^{k}+B_{0}$.
@@
    It is easy to see that $\bigcap\limits_{i=0}^{\infty}B_{i}^{\prime}=m^{k}\bigcap\limits_{i=0}^{\infty}B_{i}=m^{k}x$, where $x$ is the outcome of some game with the starting ball $B_{0}$, i.e. $x\notin S$, therefore $m^{k}x\notin S$ and Bob wins.
@@
    The strategy and the argument is analogous; we just take $A_{i}=A_{i}^{\prime}-m^{k}$ and $B_{i}^{\prime}=B_{i}+m^{k}$ and use the fact that $x\in S$ if and only if $m^{k}+x\in S$.
@@
Fix $m\in\mathbb{N}$. Let $S\subseteq\mathbb{R}$ be some set with a property that $m^{k}S=m^{k}+S=S$ for any integer $k$. Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $S$. Then for any $L>0$ and any $\varepsilon>0$ there exists a winning strategy in this game for Bob starting with $B_{0}$ with
@@
Fix $m\in\mathbb{N}$. Let $S\subseteq\mathbb{R}$ be some set with a property that $m^{k}S=m^{k}+S=S$ for any integer $k$. Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $S$. Then for any $\beta^{\prime}<\beta,\,\alpha^{\prime}=\frac{\alpha\beta}{\beta^{\prime}}$, and any closed ball $T$ Bob can win $(\alpha^{\prime},\beta^{\prime})$-game starting with any ball $T$ as his initial ball.
@@
Lemma 3.7 makes the game ”symmetric”: now we can assume that both Alice and Bob can not choose their starting balls. This symmetry makes it possible to utilize symmetric strategies:
@@
Let $m\in\mathbb{N}$ and $S,Q\subseteq\mathbb{R}$ be some sets with the following properties:
  • $m^{k}S=m^{k}+S=S$ for any integer $k$.
  • $S\sqcup(-S)\sqcup Q=\mathbb{R}$.
@@
    If $\alpha\geq\beta$, then $(\alpha,\beta)$ is a non-winning point for $S$;
@@
    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then $(\alpha,\beta)$ is a non-losing point for $S\sqcup Q$.
@@
    Suppose $\alpha\geq\beta$ and $S$ is $(\alpha,\beta)$-winning. In this case Alice has a positional winning strategy described by the function $A(I)$ (where $I\subset\mathbb{R}$ is a closed interval) for this game. Consider $(\beta,\alpha)$-game on the set $S$ and let Bob play the following strategy:
@@
    Suppose $x=\bigcap\limits_{i=0}^{\infty}B_{i}$ is the outcome of the game. Then $-x=\bigcap\limits_{i=0}^{\infty}A(-A_{n})$ is the outcome of the $(\alpha,\beta)$-game where Bob plays the balls $[-1;0];-A_{0},-A_{1},\ldots$ and Alice plays using her winning strategy $A(I)$. Therefore, $-x\in S$, which means $-x\notin S$ and Bob wins.
    It means that the point $(\alpha,\beta)$ is winning and the point $(\beta,\alpha)$ is losing for the set $S$, which contradicts Proposition 3.1. Thus, the point $(\alpha,\beta)$ is non-winning.
@@
    The idea of the proof is the same as in (a). Suppose $S\sqcup Q$ is $(\alpha,\beta)$-losing; Take $\varepsilon>0$ under the condition $(1+\varepsilon)^{2}<\frac{\beta}{\alpha}$. Then by Lemma 3.7 for any closed ball $T$ Bob has a winning strategy for $\left(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)}\right)$-game starting from $T$. We use this statement to construct a $\left(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon)\right)$-winning strategy for Alice.
    Suppose Bob starts some $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$-game with the ball $B_{0}$. Then Alice picks some ball $A_{0}\subseteq B_{0}$ without any restrictions. If Bob plays the ball $B_{n}$ on $n$-th step, Alice responds by playing $A_{n}$ where $-A_{n}$ is the ball Bob would play in $(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)})$-game which started from $-A_{0}$ after the balls $-A_{0},-B_{1},\ldots,-B_{n}$ (in this game $-A_{i}$ are Bob’s balls and $-B_{i}$ are Alice’s). Let $x$ be the outcome of this $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$-game; then $-x$ is the outcome of $(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)})$-game Bob played with his winning strategy, i.e. $-x\in(-S)$ and $x\in S\subseteq S\sqcup Q$. It shows that $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$ is a winning point for $S\sqcup Q$.
    But one has $\alpha<\alpha(1+\varepsilon)<\frac{\beta}{(1+\varepsilon)}$, which by Proposition 3.1 contradicts our initial assumption. Therefore $S\sqcup Q$ is $(\alpha,\beta)$-non-losing.
Remark 3.9.
In general, it’s not guaranteed that the game on $S$ as in Lemma 3.8 is determined, so we can only prove the weaker statements about $(\alpha,\beta)$ being non-winning or non-losing. Indeed: by ergodicity of the multiplication by $m$ map, one can see that such a $S$ has to have measure zero or be non-measurable. In the latter case, such a $S$ is not compatible with the Axiom of determinacy. It is not clear if there exists such a non-measurable $S$ on which the game is determined; we only were able to construct the examples of such $S$ using the axiom of choice.
We finish this section with an almost obvious technical statement that we need to prove Proposition 4.3.
@@
Suppose $E$ is an $(\alpha,\beta)$-winning set and for some $\xi\in\mathbb{R}$ one has $E+\xi\subseteq S$. Then the set $S$ is $(\alpha,\beta)$-winning.